If tan a= (sin a- cos a) ÷ (sin a + cos a) , then sin a+ cos a=?

Grade:10

1 Answers

Arun
6 years ago
Dear Pooja
 
tan A = sin - sin(90-A)/ (sinA + sin(90-a))
tan A = 2 cos 45 sin (45 +A)/ 2 sin 45 cos (45+A)
tan A = tan (45 +A)
A = 45 +A
This is never true.
 
Hence I think this question is wrong.